(Previous) Prof. Marcia Grail is experimenting with energy and electric charge to better create a new doomsday weapon (and take over the world). She gathers a charge of 1.11 x 10^-10 C and places it in the center of a large empty spherical vacuum chamber of radius 250m (the place she fixes this immovable charge can be thought of as the origin of coordinates). The potential at a point 1m away in the positive y-direction is 1 V. What is the electric potential 1m away in the positive x-direction? What is the direction of the electric field at 1m away in the negative y-direction? What is the direction of the electric field at 1m away in the positive x-direction? (Don't need to answer)

(Please help answer:) Satisfied with her previous experiment in the vacuum chamber, Prof. Marcia Grail needs to store a great deal of energy in order to launch her next attempt at world domination. She decides to store this energy in a gigantic parallel plate capacitor, with the area of the plates fixed, and the distance between the plates minutely tunable. This capacitor is placed in the vacuum chamber from the previous, and the atmosphere removed. If the capacitance is 0.01 Farad, and is meant to store an energy of 20 GJ, then what voltage would need to be applied? How much charge does this correspond to?
Answer #1

capacitance = C = 0.01 F

energy = U = 20 X 10^9 joules

a) U = 1/2 * C * V^2

    20 X 10^9 = 0.5 * 0.01 * V^2

Voltage = V = 2 X 10^6 Volts

b) c = q / v

charge = q = c v

                 q = 0.01 * 2 x 10^6 = 0.02 x 10^6 C